OpenStudy (mollynicole2015):

Solve for k. 3(k – 10) = 9

OpenStudy (alexandervonhumboldt2):

first distribute: 3(k-10)=9 3k-30=9 now add 30 to both sides: 3k-30+30=9+30 3k=39 now divide both sides by 3: 3k/3=39/3 k=39/3???

OpenStudy (triciaal):

you can first divide each side by 3 isolate the term with k to one side simplify

OpenStudy (alexandervonhumboldt2):

yeah there are 2 methods

OpenStudy (alexandervonhumboldt2):

it would be: 3*(k-10)=9 3*(k-10)/3=9*3 k-10=3 add 10: k-10+10=3+10 k=3+10=simplify

OpenStudy (anonymous):

Divide each side by 3 k-10 = 3 Move -10 to the RHS and simplify. k = 13

OpenStudy (vickyisthesmartone123):

Try mathway.com it help you with anything trust me type in your equation and it gives you the right answer!

OpenStudy (mollynicole2015):

Thank you! That helped! ;)